the antecedent and consequent of a ratio differ by 40 and is equal to 2/7. find the ratio.
Answer by drk(1908)

We have a fraction of the following type
N/D, where D NOT = 0.
Based on the given information,
D - N = 40
and
N/D = 2/7.
This is a reduced number.
D = 40 + N
by substitution, we get
N / (40+N) = 2/7
cross multiply to get
7N = 80 + 2N
solve for N to get
5N = 80
N = 16
Since N = 16, D = 56
Our fraction is 16/56.